# Turán number of complete multipartite graphs in multipartite graphs

### v1

@@ 3. $(K3)$
$\leq e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})+\sum_{j\neq{i_{0}}}e_{G}(Z_{j}^{i_{0}})+e_{T}(W_{i_{0}}^{\prime})+|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i_{0}}^{\prime}|-(3/4)\varepsilon n$
$\leq\alpha_{i_{0}}+\sum_{j\neq{i_{0}}}e_{G}(Z_{j}^{i_{0}})-(3/4)\varepsilon n.$ (4.20)
Combining (4.5) and (4.20) gives $\sum_{i\in[r]}e_{G}(\tilde{U}_{i}\setminus Z^{i}_{i})\leq\sum_{i\in[r]}\alpha_{i}+\sum_{i,j\neq i}e_{G}(Z_{j}^{i})-(3/4)\varepsilon n$. Since $\sum_{i,j\neq i}e_{G}(Z_{j}^{i})\leq a^{2}r^{2}t\sqrt{\gamma}n$ by (4.17) and $\gamma\ll\varepsilon$, it follows that
$\sum_{i\in[r]}e_{G}(\tilde{U}_{i}\setminus Z^{i}_{i})\leq\sum_{i\in[r]}\alpha_{i}-\varepsilon n/2$ (4.21)
Recall that $f_{1}\leq z_{t}(n)$ by (4.16). We next bound $f_{i}$ for $i\geq 2$, noting that
$f_{i}=e\left(Z_{i}^{i},G[Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})]\right)+e_{G}\left(Z_{i}^{i},(\tilde{Z}_{i}\setminus Z_{i})\cup(W^{\prime}_{i}\cap V_{q_{1}})\right).$
By Claim 4.9 (2), $G[Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})]$ has the maximum degree at most $t-1$, and thus,
$e\left(Z_{i}^{i},G[Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}})]\right)\leq e_{G}(Z_{i}^{i}\cup(W_{i}^{\prime}\setminus V_{q_{1}}))\leq\frac{t-1}{2}(|Z_{i}^{i}|+|W_{i}^{\prime}\setminus V_{q_{1}}|).$
By (4.12), $e\left(Z_{i}^{i},(\tilde{Z}_{i}\setminus Z_{i})\cup(W^{\prime}_{i}\cap V_{q_{1}})\right)\leq(t-1)\left(|\tilde{Z}_{i}\setminus Z_{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|\right)$. Putting these together,
$f_{i}\leq\frac{t-1}{2}(|Z_{i}^{i}|+|W^{\prime}_{i}\setminus V_{q_{1}}|)+(t-1)\left(|\tilde{Z}_{i}\setminus Z_{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|\right)=\frac{t-1}{2}\left(|\tilde{U}_{i}|+|\tilde{Z}_{i}\setminus Z_{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|\right).$
@@
By $(P3)$ and (4.19), we have $|Z^{1}_{1}|\geq(1-\sqrt{\gamma})an$, $|W^{\prime}_{1}|\geq(1-\gamma)n$, and thus, $|\tilde{U}_{1}|\geq(a+1)n(1-\sqrt{\gamma})$. By $(P3)$ and (4.19), we also have $|\tilde{Z}_{i}\setminus Z_{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|\leq(r-1)\sqrt{\gamma}n+\gamma n\leq r\sqrt{\gamma}n$. It follows that
$\sum_{i=2}^{r}f_{i}\leq\frac{t-1}{2}\left(kn-(a+1)n(1-\sqrt{\gamma})+(r-1)r\sqrt{\gamma}n\right)\leq(k-a-1)n\frac{t-1}{2}+\sqrt[3]{\gamma}n.$ (4.22)
@@
Since $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$, we rewrite (4.6) as $e(G)=e(G^{\prime})+\sum_{i\in[r]}f_{i}+e_{G}(\tilde{U}_{i}\setminus Z^{i}_{i})$. By Claim 4.4, it follows that $e(G)\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(f_{i}+\beta_{i}-\alpha_{i}+e_{G}(\tilde{U}_{i}\setminus Z^{i}_{i}))$. Recall that $\sum_{i=1}^{r}\beta_{i}\leq 2(r-1)rC_{0}\sqrt{\gamma}n$ by (4.8). Together with (4.21) and(4.22), we derive that
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(k-a-1)n\frac{t-1}{2}+\sqrt[3]{\gamma}n+2(r-1)rC_{0}\sqrt{\gamma}n-\varepsilon n/2<g_{2}(n,r,k,t),$
a contradiction. ∎
Next we upper bound $e(\tilde{U}_{i})$ and separate the discussions depending on whether $|W_{i}|\geq 2\varepsilon n$ or not. For $i\in[2,r]$, let $U_{i}^{*}:=\tilde{U}_{i}\setminus(Z_{1}\cup V_{q_{1}})$.
Claim 4.13.
For $i\in[2,r]$, we have
$e_{G}(\tilde{U}_{i})\leq\begin{cases}(t-1)\min\{|Z_{i}^{i}|,|W_{i}^{\prime}|\}\hfill\text{ if }a=1,\text{ and }|W_{i}|\geq 2\varepsilon n\\ |\tilde{U}_{i}|(t-1)/2\hfill\text{ if }a\geq 2\text{ and }|W_{i}|\geq 2\varepsilon n\\ (t-1)(|\tilde{Z}_{i}^{\prime}\setminus Z_{i}|+|W_{i}^{\prime}|)+\alpha_{i}\hfill\text{ if }a=1,\text{ and }|W_{i}|<2\varepsilon n\\ \left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+t^{\prime}+\alpha_{i}+a^{2}t(|Z_{1}^{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|)\hfill\text{ if }a\geq 2\text{ and}|W_{i}|<2\varepsilon n,\end{cases}$
where $t^{\prime}:=(t-1)^{2}/(16(a-1))$. Moreover, when $|W_{i}|<2\varepsilon n$, $a\geq 2$ and $t<a$, we have $e(\tilde{U}_{i})\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+\alpha_{i}+a^{2}t(|Z_{1}^{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|)$.
Proof.
Assume $|W_{i}|\geq 2\varepsilon n$. Since $|W^{\prime}_{1}\geq(1-\gamma)n$, we know that $q_{1}\neq q_{i}$. It follows from Claim 4.12 that $\tilde{U}_{i}\subseteq Z_{i}^{i}\cup(W^{\prime}_{i}\cap V_{q_{i}})$. By Claim 4.9 (2), $\tilde{U}_{i}$ is $K_{1,t}$-free and thus, $e_{G}(\tilde{U}_{i})\leq|\tilde{U}_{i}|(t-1)/2$. For $a=1$, since both $Z_{i}^{i}$ and $W_{i}^{\prime}\subseteq V_{q_{i}}$ are independent sets, it follows that $e_{G}(\tilde{U}_{i})\leq(t-1)\min\{|Z_{i}^{i}|,|W_{i}^{\prime}|\}$.
Now we consider the case when $|W_{i}|<2\varepsilon n$. Recall that
$e_{G}(\tilde{U}_{i})=f_{i}+e_{G}((Z^{\prime}_{i}\setminus Z_{i})\cup W^{\prime}_{i}).$
When $a=1$, we have $e(Z_{j}^{i})=0$ for all $j\in[r]$. Thus $f_{i}\leq(t-1)(|\tilde{Z}_{i}^{\prime}\setminus Z_{i}|+|W_{i}^{\prime}|)$ by (4.12) and $e_{G}((Z^{\prime}_{i}\setminus Z_{i})\cup W^{\prime}_{i})\leq\alpha_{i}$ by (4.5). As a result, $e_{G}(\tilde{U}_{i})\leq(t-1)(|\tilde{Z}_{i}^{\prime}\setminus Z_{i}|+|W_{i}^{\prime}|)+\alpha_{i}$ as desired.
Now assume $a\geq 2$. Let $Z^{*}_{i}:=(\tilde{Z}_{i}\setminus Z_{i})\setminus Z_{1}^{i}$, and $W^{*}_{i}:=W_{i}^{\prime}\setminus V_{q_{1}}$. For every $j\in[r]\setminus\{1,i\}$, consider the family ${\mathcal{F}}_{j}:=\{Z_{j}^{i}\cap V_{(j-1)a+1},\dots,Z_{j}^{i}\cap V_{ja},W^{*}_{i}\cap V_{q_{j}}\}$. Let $B_{j}$ be the smallest set in ${\mathcal{F}}_{j}$ and $A_{j}$ be an arbitrary set in ${\mathcal{F}}_{j}\setminus\{B_{j},W^{*}_{i}\cap V_{q_{j}}\}$. Further, define $A:=\bigcup_{j\in[r]\setminus\{1,i\}}A_{j}$ and $B:=\bigcup_{j\in[r]\setminus\{1,i\}}B_{j}$. By Claim 4.9, both $U^{*}_{i}\setminus A$ and $U^{*}_{i}\setminus B$ are $K_{1,t}$-free. This implies that $e_{G}(U^{*}_{i}\setminus B)\leq\left\lfloor|U^{*}_{i}\setminus B|\frac{t-1}{2}\right\rfloor$ and $e(B,G[U^{*}_{i}\setminus A])\leq|B|(t-1)$. Together with $e_{G}(A,B)\leq|A||B|$, we obtain that
$e_{G}(U^{*}_{i})\leq\left\lfloor|U^{*}_{i}\setminus B|\frac{t-1}{2}\right\rfloor+|B|(t-1)+|A||B|=\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+|B|\left(\frac{t-1}{2}+|A|\right).$
Since, by definition $|B|\leq(|Z^{*}_{i}|-|A|)/(a-1)=:x$, unless $\frac{t-1}{2}+|A|-|Z^{*}_{i}|<0$, we have
$|B|\left(\frac{t-1}{2}+|A|-|Z^{*}_{i}|\right)\leq x\left(\frac{t-1}{2}-(a-1)x\right)\leq\frac{(t-1)^{2}}{16(a-1)}=t^{\prime}.$
Trivially $|A|\leq|Z_{i}^{*}|$ and $|B|\leq|W^{*}_{i}|$. Altogether, we get
$e_{G}(U^{*}_{i})\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+t^{\prime}+|Z^{*}_{i}||W^{*}_{i}|.$ (4.23)
Let $X_{i}:=Z_{1}^{i}\cup(W_{i}^{\prime}\cap V_{q_{1}})$ and we bound $e(X_{i},G[\tilde{U}_{i}])$ as follows. Since $\tilde{U}_{i}=\tilde{Z}_{i}\cup W^{\prime}_{i}=Z_{i}^{i}\cup X_{i}\cup Z^{*}_{i}\cup W^{*}_{i}$, we have
$e(X_{i},G[\tilde{U}_{i}])$ $=e_{G}(X_{i},Z_{i}^{i})+e(Z_{1}^{i},G[Z_{1}^{i}\cup Z^{*}_{i}\cup W^{\prime}_{i}])+e(W^{\prime}_{i}\cap V_{q_{1}},G[Z^{*}\cup W^{\prime}_{i}])$
$=e_{G}(X_{i},Z_{i}^{i})+e_{G}(Z_{1}^{i})+e_{G}(Z_{1}^{i},Z^{*}_{i}\cup W^{\prime}_{i})+e_{G}(W^{\prime}_{i}\cap V_{q_{1}},Z^{*}_{i}\cup W^{*}_{i}).$
Note that $e_{G}(Z_{1}^{i},Z^{*}_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})$, $e_{G}(W^{\prime}_{i}\cap V_{q_{1}},W^{*}_{i})\leq e_{T}(W^{\prime}_{i})$, and
$e_{G}(Z^{i}_{1},W^{\prime}_{i})+e(W^{\prime}_{1}\cap V_{q_{1}},Z^{*}_{i})\leq|Z^{i}_{1}||W^{\prime}_{i}|+|W^{\prime}_{1}\cap V_{q_{1}}||Z^{*}_{i}|=|\tilde{Z}_{i}\setminus Z_{i}||W^{\prime}_{i}|-|Z^{*}||W^{*}|.$
Recall that $\alpha_{i}=|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|+e_{T}(W_{i}^{\prime})+e_{T}(\tilde{Z}_{i}\setminus Z_{i})$. It follows that
$e_{G}(Z_{1}^{i},Z^{*}_{i}\cup W^{\prime}_{i})+e_{G}(W^{\prime}_{i}\cap V_{q_{1}},W^{*}_{i}\cup Z^{*}_{i})$ $\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})+e_{T}(W^{\prime}_{i})+|\tilde{Z}_{i}\setminus Z_{i}||W^{\prime}_{i}|-|Z^{*}||W^{*}|$
$=\alpha_{i}-|Z^{*}||W^{*}|.$
Applying $e_{G}(X_{i},Z_{i}^{i})\leq(t-1)|X_{i}|$ from (4.12) and $e_{G}(Z_{1}^{i})\leq\binom{a}{2}(t-1)|Z_{1}^{i}|$ from Claim 4.10, we derive that
$e(X_{i},G[\tilde{U}_{i}])$ $\leq(t-1)|X_{i}|+\binom{a}{2}(t-1)|Z_{1}^{i}|+\alpha_{i}-|Z^{*}_{i}||W^{*}_{i}|.$ (4.24)
Together with (4.23), this gives the desired upper bound
$e_{G}(\tilde{U}_{i})$ $\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+t^{\prime}+(t-1)|X_{i}|+\binom{a}{2}(t-1)|Z_{1}^{i}|+\alpha_{i}$
$\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+t^{\prime}+a^{2}t|X_{i}|+\alpha_{i}.$
At last, when $a\geq 2$ and $t<a$, we know that $U^{*}_{i}$ is $K_{1,t}$-free (Claim 4.9 (3)) and thus $e_{G}(U^{*}_{i})\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor$. Together with (4.24), this gives $e_{G}(\tilde{U}_{i})\leq\left\lfloor|U^{*}_{i}|\frac{t-1}{2}\right\rfloor+\alpha_{i}+a^{2}t(|Z_{1}^{i}|+|W_{i}^{\prime}\cap V_{q_{1}}|)$. ∎
For $i\in[r]$, if $|W_{i}|\geq 2\varepsilon n$, then $\tilde{Z}_{i}\setminus Z_{i}=\emptyset$ by Claim 4.12, and therefore, $\beta_{i}=0$. If $|W_{i}|<2\varepsilon n$, then $Z_{i}\setminus\tilde{Z}_{i}=\emptyset$ by $(Q2)$; for any $j\in L\setminus\{i\}$, we have $\tilde{Z}_{j}\setminus Z_{j}=\emptyset$ and $W^{\prime}_{j}\subseteq V_{q_{i}}$ again by Claim 4.12. Hence $\beta_{i}=\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)\leq 0$ because $|W^{\prime}_{j}|\leq n$. Together with (4.16), for $i\in[r]$,
$\beta_{i}=\begin{cases}0&\text{ if $|W_{i}|\geq 2\varepsilon n$ (including $i=1$)},\\ \sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|W_{j}^{\prime}|-n)\leq 0&\text{ otherwise}.\end{cases}$ (4.25)
We are ready to finish our proof. Write $z_{i}:=|\tilde{Z}_{i}|$ and $w_{i}:=|W_{i}^{\prime}|$ for $i\in[r]$. Further, write $n_{l}:=|Z_{1}^{1}\cap V_{l}|$ for $\in[a]$. Then $z_{1}=\sum_{l\in[a]}n_{l}$. By (4.16), we have $e_{G}(\tilde{U}_{1})=f_{1}\leq z_{t}(n_{1},\dots,n_{a},w_{1})$.
We separate the cases when $a\geq 2$ and when $a=1$.
The case when $a\geq 2$
Note that there are at most $r-b$ indices $i\in[r]$ such that $|W_{i}|<2\varepsilon n$. By Claim 4.13, and using that $\sum_{i=2}^{r}|U^{*}_{i}|=(k-a-1)n$, we have
$\sum_{i\in[r]}e_{G}(\tilde{U}_{i})\leq f_{1}+(an-z_{1}+n-w_{1})a^{2}t+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor+(r-b)t^{\prime}+\sum_{i:|W_{i}|<2\varepsilon n}\alpha_{i}$
Since $e(G)=e(G^{\prime})+\sum_{i\in[r]}e_{G}(\tilde{U}_{i})$, it follows that
$e(G)\leq f_{1}+(an-z_{1}+n-w_{1})a^{2}t+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor+(r-b)t^{\prime}+e(G^{\prime})+\sum_{i:|W_{i}|<2\varepsilon n}\alpha_{i}.$
Claim 4.4 and (4.25) together imply that
$e(G^{\prime})\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(\beta_{i}-\alpha_{i})\leq t_{r}(k)n^{2}-\sum_{i:|W_{i}|<2\varepsilon n}\alpha_{i}.$
Together with $f_{1}\leq z_{t}(n_{1},\dots,n_{a},w_{1})$, this gives
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n_{1},\dots,n_{a},w_{1})+(an-z_{1}+n-w_{1})a^{2}t+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor+(r-b)t^{\prime}.$
If $an-z_{1}+n-w_{1}>0$, then by applying ($E3$) repeatedly, we have
$z_{t}(n_{1},\dots,n_{a},w_{1})+(an-z_{1}+n-w_{1})a^{2}t+(r-b)t^{\prime}\leq z_{t}(n).$
This implies that $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor$. Otherwise $an+n=z_{1}+w_{1}$, which implies that
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor+(r-b)t^{\prime}.$
This is the desired bound for $t\geq a$. When $t<a$, the bound for $e_{G}(\tilde{U}_{i})$ given by Claim 4.13 contains no $t^{\prime}$ term. We thus obtain $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+\left\lfloor\frac{t-1}{2}(k-a-1)n\right\rfloor$, regardless whether $an-z_{1}+n-w_{1}>0$.
Now consider the case $a\geq 2$ and $b=1$. If $an-z_{1}+n-w_{1}>0$, then we choose a large constant $C^{*}>0$ such that
$z_{t}(n_{1},\dots,n_{a},w_{1})+(an-z_{1}+n-w_{1})a^{2}t+(r-b)t^{\prime}\leq z_{t}(n)-C^{*}$
and thus $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+\frac{t-1}{2}(k-a-1)n-C^{*}<g_{2}(n,r,k,t)$. Otherwise $an+n=z_{1}+w_{1}$. Since $b=1$, this forces that, for all $i\geq 2$, we have $W_{i}^{\prime}=\emptyset$ and, by $(Q2)$,$\tilde{Z}_{i}=Z_{i}$. Now, by Claim 4.9, we have $\sum_{i=2}^{r}e_{G}(\tilde{U}_{i})\leq(r-1)\lfloor\frac{t-1}{2}an\rfloor$ and consequently, $e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(r-1)\lfloor\frac{t-1}{2}an\rfloor$, as desired.
The case when $a=1$
@@ 2. $(R2)$
    If $i\in L_{2}$, then $\tilde{Z}_{i}=Z_{i}^{i}=V_{i}$. Otherwise $\tilde{Z}_{i}\neq Z_{i}^{i}$, then $|Z_{i}^{i}|<n$ and $Z_{i}^{j}\neq\emptyset$ for some $j\neq i$. By $(Q2)$ and (4.12), we have $\tilde{Z}_{i}=Z_{i}^{i}$, a contradiction.
@@ 3. $(R3)$
    If $i\in L_{3}$, then $\tilde{Z}_{i}\not\subseteq Z_{i}$. By Claim 4.12, we have $|W_{i}|<2\varepsilon n$, which implies that $Z_{i}^{j}=\emptyset$ for $j\neq i$ by $(Q2)$. Thus, $Z_{i}^{i}=V_{i}\subsetneq\tilde{Z}_{i}$.
By $(R1)$, for every $i\in L_{1}$, $G[\tilde{U}_{i}]$ is a bipartite graph of maximum degree at most $t-1$, and thus $e_{G}(\tilde{U}_{i})\leq(t-1)\min\{z_{i},w_{i}\}$. For $i\in L_{2}$, since $\tilde{Z}_{i}=V_{i}$ is an independent set, we have $e_{G}(\tilde{U}_{i})\leq(t-1)w_{i}+e_{G}(W_{i}^{\prime})\leq(t-1)w_{i}+\alpha_{i}$. For $i\in L_{3}$, using (4.5) and the independence of $Z_{j}^{i}$ for all $j\in[r]$, we have
$e_{G}(\tilde{U}_{i})$ $=e_{G}((\tilde{Z}_{i}\setminus V_{i})\cup W_{i}^{\prime})+e_{G}((\tilde{Z}_{i}\setminus V_{i})\cup W_{i}^{\prime},Z_{i}^{i})$
$\leq\alpha_{i}+(t-1)(|\tilde{Z}_{i}\setminus V_{i}|+|W_{i}^{\prime}|)=\alpha_{i}+(t-1)(z_{i}-n+w_{i}).$
Using $\sum_{i\in L_{1}\cup L_{3}}(z_{i}-n)=n-z_{1}$ and $\sum_{i=2}^{r}w_{i}=bn-w_{1}$, we derive that
$\sum_{i\in L_{1}}\min\{z_{i},w_{i}\}+\sum_{i\in L_{2}}w_{i}+\sum_{i\in L_{3}}(z_{i}-n+w_{i})$
$=$ $bn-w_{1}+\sum_{i\in L_{1}}(\min\{z_{i},w_{i}\}-w_{i})+\sum_{i\in L_{3}}(z_{i}-n)$
$=$ $bn-w_{1}+\sum_{i\in L_{1}}(z_{i}-n+\min\{n-w_{i},n-z_{i}\})+\sum_{i\in L_{3}}(z_{i}-n)$
$=$ $bn-w_{1}+n-z_{1}+\sum_{i\in L_{1}}\min\{n-w_{i},n-z_{i}\}.$

### v2

@@ 3. $(K3)$
$\leq e_{T}(\tilde{Z}_{i_{0}}\setminus Z_{i_{0}})+e_{T}(W_{i_{0}}^{\prime})+|\tilde{Z}_{i_{0}}\setminus Z_{i_{0}}||W_{i_{0}}^{\prime}|-(1/2)\varepsilon n$
$\leq\alpha_{i_{0}}-(1/2)\varepsilon n.$ (4.17)
Combining (4.5) and (4.17) gives
$\sum_{i\in[r]}e_{G}(\tilde{U}_{i}\setminus Z_{i})\leq\sum_{i\in[r]}\alpha_{i}-(1/2)\varepsilon n$ (4.18)
@@
Recall that $f_{i}\leq(t-1)|\tilde{U}_{i}\setminus Z_{i}|$ ($i\geq 2$) by (4.11), $|W^{\prime}_{1}|\geq(1-\gamma)n$ by (4.16), and $|Z_{i}^{i}|\geq(1-\sqrt{\gamma})n$ by $(P3)$. Therefore, as $\sum_{i=1}^{r}|\tilde{U}_{i}\setminus Z_{i}|\leq(b-1)n+r\sqrt{\gamma}n+{\gamma}n$, we obtain
$\sum_{i=2}^{r}f_{i}\leq(t-1)\left((b-1)n+r\sqrt{\gamma}n+{\gamma}n\right)\leq(t-1)(b-1)n+\sqrt[3]{\gamma}n.$ (4.19)
@@
Since $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$, (4.7) becomes $e(G)\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(f_{i}+\beta_{i}+e_{G}(\tilde{U}_{i}\setminus Z_{i})-\alpha_{i})$. Recall that $\sum_{i=1}^{r}\beta_{i}\leq 2r^{2}C_{0}\sqrt{\gamma}n$ by (4.8). Together with (4.18) and (4.19), we derive that
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(b-1)n+\sqrt[3]{\gamma}n+2r^{2}C_{0}\sqrt{\gamma}n-\varepsilon n/2<g(n,r,k,t),$
as $\gamma\ll\varepsilon$. This is a contradiction. ∎
@@ 2. $(R2)$
    If $i\in L_{2}$, then $\tilde{Z}_{i}=Z_{i}^{i}=V_{i}$. Indeed, otherwise $\tilde{Z}_{i}\neq Z_{i}^{i}$, then $|Z_{i}^{i}|<n$ and $Z_{i}^{j}\neq\emptyset$ for some $j\neq i$. By $(Q2)$ and Claim 4.10, we have $\tilde{Z}_{i}=Z_{i}^{i}$, a contradiction.
@@ 3. $(R3)$
    If $i\in L_{3}$, then $\tilde{Z}_{i}\not\subseteq Z_{i}$ (otherwise $|\tilde{Z}_{i}|\leq n$). By Claim 4.10, we have $|W_{i}|<2\varepsilon n$, which implies that $Z_{i}^{j}=\emptyset$ for $j\neq i$ by $(Q2)$. Thus, $Z_{i}^{i}=Z_{i}=V_{i}\subsetneq\tilde{Z}_{i}$.
Now we derive our final bound on $e(G)$. Write $z_{i}:=|\tilde{Z}_{i}|$ and $w_{i}:=|W_{i}^{\prime}|$ for $i\in[r]$.
By Claim 4.3 and the fact that $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$, we have
$e(G)=e(G^{\prime})+\sum_{i\in[r]}e_{G}(\tilde{U}_{i})\leq t_{r}(k)n^{2}+\sum_{i\in[r]}\left(\beta_{i}-\alpha_{i}+e_{G}(\tilde{U}_{i})\right).$
Moreover, as $a=1$, (4.5) becomes $e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})\leq\alpha_{i}$. It follows that $e_{G}(\tilde{U}_{i})=f_{i}+e_{G}(\tilde{U}_{i}\setminus Z_{i})\leq f_{i}+\alpha_{i}$. For $i\in\{1\}\cup L_{1}\cup L_{2}$, we simply use $f_{i}$ as the upper bound and thus we get
[…]
Additional work is needed for $i\in L_{3}$. We let $\lambda=\max\{0,|L_{1}|-|L_{3}|\}$ and for $i\in L_{3}$, let $\lambda_{i}$ be the number of indices $j\in L_{1}$ such that $Z_{j}^{i}\neq\emptyset$. By $(R1)$$(R3)$, we know that if $Z_{j}^{i}\neq\emptyset$, then $i\in L_{3}$ and $j\in L_{1}$. This implies that $\lambda_{i}\geq 1$ for every $i\in L_{3}$, and $\sum_{i\in L_{3}}\lambda_{i}\geq|L_{1}|$, yielding that
$\sum_{i\in L_{3}}(\lambda_{i}-1)\geq\lambda.$ (4.20)
Recall that $G[tZ_{i}\setminus Z_{1}^{i}]$ is $K_{1,t}$-free by Claim 4.9 (2). Since $Z_{i}^{i}=Z_{i}$ is an independent set, it follows that $e_{G}(\tilde{Z}_{i}\setminus Z_{1}^{i})\leq(t-1)|\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}^{i})|$. Together with (4.11), this gives
$e_{G}(\tilde{Z}_{i}\setminus Z_{1}^{i})+e_{G}(Z_{1}^{i},Z_{i})\leq(t-1)|\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}^{i})|+(t-1)|Z_{1}^{i}|=(t-1)(z_{i}-n).$
Therefore, for $i\in L_{3}$, writing $\varrho_{i}:=e_{T}(Z_{1}^{i},\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}^{i}))$, we have
$e_{G}(\tilde{Z}_{i})=e_{G}(\tilde{Z}_{i}\setminus Z_{1}^{i})+e_{G}(Z_{1}^{i},Z_{i})+e_{G}(Z_{1}^{i},\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}^{i})\leq(t-1)(z_{i}-n)+\varrho_{i}.$
Moreover, the definition of $\alpha_{i}$ implies that
$e_{G}(W_{i}^{\prime})+e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})-\alpha_{i}$ $\leq e_{T}(W_{i}^{\prime})+|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|-\alpha_{i}=-e_{T}(\tilde{Z}_{i}\setminus Z_{i})$
$=-\varrho_{i}-e_{T}(\tilde{Z}_{i}\setminus(Z_{i}\cup Z_{1}))$
$\leq-\varrho_{i}-\binom{\lambda_{i}}{2}\leq-\varrho_{i}+1-\lambda_{i}.$
Finally, by (4.11), we have $e_{G}(Z_{i},W_{i}^{\prime})\leq(t-1)w_{i}$ for $i\in L_{3}$. Combining all these inequalities together, we obtain that, for $i\in L_{3}$,
$e_{G}(\tilde{U}_{i})-\alpha_{i}$ $=e_{G}(\tilde{Z}_{i})+e_{G}(Z_{i},W_{i}^{\prime})+e_{G}(W_{i}^{\prime})+e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})-\alpha_{i}$
$\leq(t-1)(z_{i}-n+w_{i})+(1-\lambda_{i}).$
It follows that $\sum_{i\in L_{3}}(e_{G}(\tilde{U}_{i})-\alpha_{i})\leq\sum_{i\in L_{3}}(t-1)(z_{i}-n+w_{i})-\lambda$ by using (4.20).
Using $\sum_{i=2}^{r}(z_{i}-n)=n-z_{1}$ and $\sum_{i=2}^{r}w_{i}=bn-w_{1}$, we derive that
$\sum_{i\in L_{1}}\min\{z_{i},w_{i}\}+\sum_{i\in L_{2}\cup L_{3}}(z_{i}-n+w_{i})=$ $\sum_{i\in L_{1}}\min\{n-w_{i},n-z_{i}\}+\sum_{i=2}^{r}(z_{i}-n+w_{i})$
$=$ $\sum_{i\in L_{1}}\min\{n-w_{i},n-z_{i}\}+bn-w_{1}+n-z_{1}.$